A post is 9 ft. from one wall of a room and 40 ft. from the wall at right angle to it. How many ft. is it from the the intersection of the two walls?
Answer by stanbon(57238)

A post is 9 ft. from one wall of a room and 40 ft. from the wall at right angle to it. How many ft. is it from the the intersection of the two walls?
------------
Let one wall be the x-axis ; Let the other wall be the y-axis.
The post is at (4,9); the interection of the walls is (0,0)
----------
Distance from (0,0) to (4,9):
d = sqrt(4^2+9^2)= sqrt(16+36)= sqrt(52) ft.
=============
Cheers,
Stan H.